# On the Robustness of Adversarial Training Against Uncertainty Attacks

### v1

@@ 5.2 Experimental Results
Here the results reflect well our analysis: the group $\diamondsuit$ comes after group $\clubsuit$; the same applies for ImageNet, where almost all models from group $\heartsuit$ come after, except I5, the transformers (group $\spadesuit$). Generally speaking - except for some outliers - we observe a slight correlation between robust accuracy and robust uncertainty: disregarding C10 and C2, group $\clubsuit$ has (on average) lower MUS and MSUS than group $\diamondsuit$, while the uncategorized models perform similarly to $\clubsuit$; at the same time - without taking I5 into account - the Transformers outperform the ResNet architectures. Nevertheless, the presence of C2 in the second place reveals that high entropic models may eventually disclose general robust properties against uncertainty attacks. It is also important to note that C10 employs Neural ODE blocks, which are known to be affected by obfuscated gradients[25, 33]. This phenomenon hinders the effectiveness of gradient-based attacks, leading to unreliable robustness evaluations. Accordingly, since our evaluation consists of gradient-based attacks, we mark the results obtained with C10 as unreliable since it may change by employing attack transferability[25, 33].
6 Conclusions, Limitations and Future Works
In this work, we provided an analysis of the robustness of adversarial-trained models against uncertainty attacks. Our experimental evaluation has empirically validated the hypothesis that instead of needing an ad hoc defense strategy for uncertainty attacks, one can rely on already trained state-of-the-art robust models. Nevertheless, it is worth underlining two limitations of our current study. Firstly, since some classifiers (e.g., Kang [25]) are robust to AutoAttack but weak against more sophisticated attacks, considering a single instance of over- and under-confidence attacks may sometimes lead to potentially unreliable results. Secondly, we limit ourselves to studying the entropy and, thus, only aleatoric uncertainty.
Accordingly, on a closing note, we aim to suggest future research directions towards (i) evaluating methods such as Kang with more sophisticated attacks (such as transfer attacks) and (ii) extending such analysis also to epistemic uncertainty, particularly for proper Bayesian models.
@@ Acknowledgments
This work has been supported by the European Union’s Horizon Europe research and innovation program under the project ELSA, grant agreement No 101070617; by Fondazione di Sardegna under the project “TrustML: Towards Machine Learning that Humans Can Trust”, CUP: F73C22001320007; by EU - NGEU National Sustainable Mobility Center (CN00000023) Italian Ministry of University and Research Decree n. 1033—17/06/2022 (Spoke 10); and by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU. This work was conducted while Emanuele Ledda, Daniele Angioni, and Giorgio Piras were enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the University of Cagliari.

### v2

@@ V-B Experimental Results
Here, the results reflect well our analysis: the group $\diamondsuit$ comes after group $\clubsuit$; the same applies for ImageNet, where almost all models from group $\heartsuit$ come after, except I5, the transformers (group $\spadesuit$). Generally speaking - except for some outliers - we observe a slight correlation between robust accuracy and robust uncertainty: disregarding C10 and C2, group $\clubsuit$ has (on average) lower MUS and MSUS than group $\diamondsuit$, while the uncategorized models perform similarly to $\clubsuit$; at the same time - without taking I5 into account - the Transformers outperform the ResNet architectures. Nevertheless, the presence of C2 in the second place reveals that high entropic models may eventually disclose general robust properties against uncertainty attacks.
It is also important to note that LABEL:tab:leaderboard reports the rankings obtained by RobustBench ordered by the best known robust accuracy. Most of these robust accuracies are obtained by using AutoAttack, which is an ensemble of gradient-based white- and black-box attacks. However, for certain models, even AutoAttack reveals to be unreliable: this is the case of C10, for which the best known robust accuracy was obtained by running a transfer attack. In fact, methods that make use of Neural ODE in their architecture like C10 are known to be affected by obfuscated gradients [25, 38]. This phenomenon hinders the effectiveness of gradient-based attacks, leading to unreliable robustness evaluations [39]. Accordingly, since our evaluation consists of gradient-based attacks, we mark the results obtained with C10 as unreliable, noting that our result may change by employing different attack strategies, e.g., attack transferability, as suggested in [25, 38].
OOD and Open-Set resultsTable V reports the results of the experiments for the OOD and OSR scenarios.
OSR OOD
Model ID $\epsilon$ AUROC AUPR-IN AUPR-OUT FPR95TPR AUROC AUPR-IN AUPR-OUT FPR95TPR
C9 clean 0.818 0.977 0.296 0.589 0.724 0.963 0.189 0.710
C9 4/255 0.741 0.966 0.175 0.695 0.681 0.956 0.151 0.770
C9 8/255 0.642 0.951 0.116 0.782 0.635 0.948 0.122 0.816
C5 clean 0.823 0.977 0.318 0.596 0.668 0.957 0.124 0.729
C5 4/255 0.750 0.967 0.191 0.687 0.635 0.951 0.112 0.772
C5 8/255 0.655 0.953 0.123 0.758 0.599 0.944 0.102 0.805
C13 clean 0.816 0.976 0.298 0.602 0.602 0.943 0.104 0.823
C13 4/255 0.744 0.966 0.182 0.699 0.572 0.937 0.096 0.858
C13 8/255 0.651 0.952 0.120 0.758 0.541 0.930 0.090 0.885
C6 clean 0.836 0.979 0.334 0.571 0.651 0.952 0.118 0.793
C6 4/255 0.767 0.969 0.204 0.666 0.621 0.946 0.109 0.828
C6 8/255 0.674 0.956 0.130 0.729 0.589 0.940 0.101 0.858
C4 clean 0.830 0.980 0.296 0.513 0.777 0.974 0.200 0.530
C4 4/255 0.764 0.971 0.186 0.607 0.762 0.972 0.184 0.549
C4 8/255 0.676 0.958 0.127 0.693 0.745 0.969 0.170 0.570
C14 clean 0.840 0.975 0.391 0.695 0.943 0.994 0.614 0.206
C14 4/255 0.587 0.915 0.116 0.989 0.930 0.993 0.542 0.244
C14 8/255 0.348 0.851 0.065 1.000 0.914 0.991 0.465 0.274
TABLE V: AUROC, AUPR and FPR95TPR computed on a mixture of ID-OSR (on the left) and ID-OOD (on the right) samples, for six Robust WideResNets, before and after the under-confidence attacks on the OOD and OSR subsets.
As we can see from the results, all the considered metrics deteriorate in the presence of an attacker. Nevertheless, even though such deterioration can vary on the base of the considered model, we notice that it does not lead to a complete break, as it happens for undefended models, where even for small perturbation budget it is possible to make the uncertainty measure on the OOD lower than any ID sample [7]; this is a clear indicator of robustness for adversarially trained models against OOD and Open Set uncertainty attacks. Quantitatively, we also notice that in almost all the cases, the adversarially trained models are able to discriminate better between open set samples compared to OOD samples. The aleatoric nature of the considered uncertainty measure can be a possible explanation of this phenomenon. Indeed, we remark that the entropy is an aleatoric measure, which is based upon the output prediction vector; notably, OOD represents an emblematic case of epistemic uncertainty, because, by definition, the considered model has no prior knowledge of the samples which are far from the training distribution. Instead, for the open set case, the lack of knowledge is less evident: even though the selected classes from CIFAR-100 were not present in the training data, some visual features may possibly be shared across classes (e.g., features from the “helicopter” class may be present in some CIFAR-10 veichle classes su as the “airplane” class).
Interestingly, we notice that C4 and C14 are more sensitive to the open set attacks compared to the remaining ones. In particular, C14 suffers from an extreme deterioration under attack in the open set case (e.g., AUROC from $0.840$ ot $0.348$), which is in line with the ID leaderboard results in LABEL:tab:leaderboard.
VI Conclusions, Limitations and Future Works
In this work, we provided an extensive analysis of the robustness of adversarial-trained models against uncertainty attacks. Our experimental evaluation conducted on image classification benchmarks has empirically validated our hypothesis that instead of needing an ad hoc defense strategy for uncertainty attacks, one can rely on already trained state-of-the-art robust models. Nevertheless, we point out some limitations of our work, which we consider promising starting points for future research. First, while our theoretical analysis reveals the interesting behaviour of adversarial trained models being underconfident by nature, this is still limited to 2-class cases and assumes the model finds a theoretical equilibrium. Nevertheless, as happens often in Machine Learning [12], it is important to notice that all the discussed theoretical results should be taken not as an ultimate theory, but rather as a compass that can guide the research process toward conclusions which are interesting from an engineering perpective, even if the theoretical assumptions do not perfectly reflect the complexity of the real world. Second, the attack setup that is shared across all considered models is not intended to be exhaustive from the adversarial perspective. For instance, we discussed in Sect. V-B how [25] needs further investigation using more advanced techniques to properly assess its robustness [25, 38, 39]. To this end, our proposed leaderboard should not be used as the ultimate benchmark, but rather as a general compass that offers valuable insights on the uncertainty robustness capabilities of adversarial trained models. Third, we prioritized our analysis to the sole aleatoric uncertainty, as previous research on adversarial robustness led to a plethora of pre-trained robust models that are deterministic by design, hence with no epistemic uncertainty available [13]. Accordingly, as future research, we intend to extend such analysis to epistemic uncertainty as well, particularly for proper Bayesian models, by employing seminal promising solutions that integrate adversarial training on top of Bayesian Neural Networks [40]. On a closing note, we believe that our study represents a first step toward understanding the robustness of standard benchmark models under adversarial perturbations targeting uncertainty quantification and their impact on uncertainty estimates. The insights obtained open up promising directions for transferring these findings to emerging areas in AI, such as generative and foundation models, as well as to safety-critical domains like autonomous driving and healthcare.
@@ Acknowledgments
This work has been supported by the European Union’s Horizon Europe research and innovation program under the projects ELSA (grant no. 101070617) and CoEvolution (grant no. 101168560); by Fondazione di Sardegna under the project “TrustML: Towards Machine Learning that Humans Can Trust”, CUP: F73C22001320007; by EU-NGEU National Sustainable Mobility Center (CN00000023) Italian MUR Decree n. 1033—17/06/2022 (Spoke 10); and by projects SERICS (PE00000014) and FAIR (PE00000013) under the NRRP MUR program funded by the EU-NGEU. This work was conducted while Emanuele Ledda, Daniele Angioni, and Giorgio Piras were enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the University of Cagliari.